The speed of a passenger train is 16mph faster then the speed of a freight train. the passenger train travels 270 miles in the same time it takes the freight train to travel 190 miles. Find the speed of each train
Answer by richwmiller(17219)

r*t=d
t=d/r
p=f+16
t=270/(f+16)
t=D/R
t=190/f
270/(f+16)=190/f
f=38 mph
p= f+16=54 mph